Planning week 9-16 May 2009

This week began with an active region appearing over the NE limb on 8 May. Another, brighter, active region followed it and I refer to them as the leading and trailing active regions, respectively. Most of the planning week was spent following the trailing active region.

Eclipse season has started so I tried to only schedule studies during the day-time period of the orbit. Note that there are about 14 orbits in a day.

The week was divided into three OP periods: OP1 was 3 days long, and OP2 and OP3 were 2 days long.

Key observations

  1. From Saturday to Tuesday I followed the trailing active region with PRY_slot_contextLITE for most of the time so there should be some nice movies. In particular, there were a few small flares during Saturday and early Sunday which might be interesting.
  2. On Wednesday I produced a 2x2 mosaic of the trailing AR using Eclipse_raster_v2, covering 360"x360". This study takes a large chunk of the EIS spectrum and so should be a nice reference data-set.
  3. The main observation of the week is the large raster scan on Friday. This should end up a contiguous 760"x400" scan covering both active regions, and the study includes many good diagnostics. XRT will be supporting with dynamics observations as well as DEM studies. SOT will be obtaining SP maps, and G and Ca images. I will be getting TRACE support with the three EUV filters, together with dynamics studies. CDS will be performing an active region mosaic.
  4. On Thursday I've scheduled a single run of Eclipse_raster_v2 at the leading footpoints of the trailing AR. The angle of the line-of-sight should be just right to get a clear view of several footpoints.
  5. There should be large slot context images of the trailing AR for most of the week. I tried to schedule at least one every orbit.

Saturday 9 May (11-24 UT)

EIT image at 13:13 UT

The trailing active region was monitored at the limb continuously from 11:20 to 05:00 on Sunday using the slot context raster PRY_slot_contextLITE (11 orbits in all). A break of 30 mins occurred during 14:30 to 15:00 due to SOT support at disk center, and an XRT synoptic was performed at 18:10.

The GOES plot reveals a number of A and B flares between 11 and 04 UT.

SOT plan (OP1)
XRT plan (OP1)

Sunday 10 May 

EIT image at 10:24 UT

The active region monitoring begun on Saturday with PRY_slot_contextLITE continued until 05:00. The next orbit was skipped due to an XRT synoptic. The monitoring continued between 07:00 and 10:00 (two orbits). 

For the next orbit (10:20), the new study Eclipse_raster_v2 was tested on the trailing active region. A single PRY_slot_context_v3 was used as context. The eclipse raster is not in the EIS archive.

The following two orbits (12:00-14:30), active region monitoring was continued with PRY_slot_contextLITE. 

The next three orbits were taken up with engineering studies and the north/east coalignment campaign. No further observations took place on Sunday.

The GOES plot reveals a number of A and B flares between 00 and 04 UT. The region is quiet until around 12 UT when a B flare occurs. The region is quiet for the rest of the day.

Monday 11 May

EIT image at 16:25 UT

Monitoring of the trailing active region continued with PRY_slot_contextLITE for 5 orbits between 01:08 and 08:44. 

The following orbit was left blank as a pointing update load was planned. Note that this update was to switch from fixed pointing to tracking. When comparing images from before and after 10:00 on Monday, there will probably be a significant spatial offset.

Active region monitoring continued for two orbits between 10:59 and 13:52, but the next 3 orbits were skipped due to data volume restrictions. Monitoring resumed for the final three orbits of the day (19:11-23:40)

study on east limb active region for 01-08 UT. Did not support HOP 73 between 08-12 UT. Ran PRY_slot_contextLITE study on east limb active region for 12-14:30 UT. Did not support HOP 118 between 14:30-18 UT.

Ran PRY_slot_contextLITE study on east limb active region for 18-23 UT. Did not support Hida campaign between 23-01 UT.

The Sun is very quiet in the GOES plot, with only an A-flare occuring at around 23:00 UT.

Tuesday 12 May (00-11 UT)

Active region monitoring with PRY_slot_contextLITE was performed on the trailing active region for only two orbits between. 01:45-04:36 UT. No other observations were performed.

The Sun is very quiet in the GOES plot.

Tuesday 12 May (11-24 UT)

EIT image at 14:15 UT

Between 10:30 and 16:03, PRY_slot_context_v3 was run continuously during the day-time periods.

Between 18:30 and 00:10, IUU_SCAN_STEPS_002 was run during four consecutive day-time periods. At the start of each the context rasterPRY_slot_context_v3 was run once. IUU_SCAN_STEPS_002 covers 68x400 arcsec^2 and has a cadence of 6mins.

The Sun is very quiet in the GOES plot.

SOT plan (OP2)
XRT plan (OP2)

Wednesday 13 May

EIT image at 12:00 UT

PRY_slot_context_v3 was run once per orbit up to 10:30, for 6 orbits in all. During 10:40 to 16:11 four orbits were used to run Eclipse_raster_v2 in a 2x2 mosaic over the trailing active region. PRY_slot_context_v3 was run as context at the start of each run. Only two of the four eclipse rasters are currently in the EIS archive.

For the remaining four orbits of the day, PRY_slot_context_v3 was run once per orbit.

Thursday 14 May (00-11 UT)

The north pole co-alignment was performed between 00:20 and 02:20 and the east limb co-alignment was performed between 04:20 and 06:20. EIS supported only during day-time periods of the orbits.

For the remaining orbits PRY_slot_context_v3 was run as context on the trailing active region once per orbit.

Thursday 14 May (11-24 UT)

EIT image at 14:48 UT

There are a group of loop footpoints on the west side of the trailing active region that might be quite interesting when the region is closer to disk center. I've thus scheduled one run of Eclipse_raster_v2 at 11:17 to observe them. There should be a co-spatial SP map fairly close in time to this observation, too. The eclipse raster at 11:17 does not exist in the EIS database.

For the rest of Thursday I scheduled one run of PRY_slot_context_v3 each orbit, following on from the previous OP period.

SOT plan (OP3)
XRT plan (OP3)

Friday 15 May

Monitoring of the trailing AR with PRY_slot_context_v3 continued until 07:30.

HOP 119 was supported between 08:34 and 11:23 using the study IUU_SLOT_148x400. Note that I ran this study with JPEG65 to reduce the data rate. The pointing was the trailing AR.

Between 11:51 and 22:54 a large raster scan was performed covering both active regions. Two orbits were spent on the leading AR, two more orbits on the region between the ARs (where there may be a filament), and three orbits on the trailing AR. I used HPW017 (30"x400"), making use of ASRC. I made sure to keep to keep the X and Y coordinates the same to ensure that all the rasters can be put together and create a continuous single raster.

The start times of the individual HPW017 rasters are given below, with comments:

11:51    Raster 1
12:07    Raster 2
12:23    Raster 3
12:39    Raster 4
13:30    Raster 5
13:46    Raster 6
14:02    Raster 7
14:18    Raster 8
15:10    Raster 9
15:26    Raster 10
15:42    Raster 11
15:58    Raster 12
16:48    Raster 13
17:04    Raster 14
17:20    Raster 15
17:36    Raster 16
18:52    Raster 17 (incorrect compensation steps)
19:08    Raster 18 (incorrect compensation steps)
20:05    Raster 19
20:21    Raster 20
20:37    Raster 21
20:53    Raster 22
21:50    Raster 23 (missing exposures? should have started at 21:44)
22:00    Raster 24
22:16    Raster 25
22:32    Raster 26

For rasters 17 and 18, I set the compensation steps incorrectly for the raster stitching. They should have been 120, but I set 60 which meant that only a 15" jump occurred between the two rasters. This only affects raster 18, though.

Fe XII image from the mosaic. The numbers near the bottom of the image denote the raster number. Note the problem with raster 18 mentioned above. Also the jump between raster 18 and raster 19 seems to be too big.

Saturday 16 May (0-11 UT)

Monitoring of the trailing AR with PRY_slot_context_v3 continued until 08:03.

HOP 119 was supported between 08:07 and 10:22 using the study IUU_SLOT_104X280_Q65. The pointing was the trailing AR.